Assume we have a closed economy, the following information for the third quarter of 2020 (assume export = import = net export = 0): 

National income (excluding net export) = 517,536,000,000 SAR.

Consumption equal 222,526,000,000 SAR.

Government expenditures equal 140,199,000,000 SAR.

Assume that the consumption tax rate is 15% (Value Added Tax) and 20% for capital (investment). Solve for the following if it is considered a closed economy:

a- Tax revenue. there are two types of tax; sales tax (VAT) and capital or investment.  First, calculate investment to find tax revenue from investment.  Add these two together.

b- Investment. 

c- Private saving.

d- Public saving and what kind of budget does Saudi Arabia has.

e- National saving.

f- On a few words; explain how the government could increase investment level using fiscal policy.
Answer #1

National income (excluding net export), Y = 517,536,000,000

Consumption , C =  222,526,000,000

Government expenditures , G = 140,199,000,000

At goods market equilibrium in a closed economy,

Y = C + I + G

=> I = Y - C - G

=> I = 517,536,000,000 - 222,526,000,000 - 140,199,000,000

=> Investment, I = 154,811,000,000

Consumer Tax = 15% of C = (15/100)*222,526,000,000 = 33,378,900,000

Investment Tax = 20% of I = (20/100)*154,811,000,000 = 30,962,200,000

Tax Revenue, T = Consumer Tax + Investment Tax = 33,378,900,000+ 30,962,200,000 = 64,341,100,000

Private Savings = Y - T - C = 517,536,000,000 -   64,341,100,000 - 222,526,000,000 = 230,668,900,000

Public Savings = T - G = 64,341,100,000 - 140,199,000,000 = - 75,857,900,000

Saudia Arabia has Budget Deficit.

National Savings = Private Savings + Public Savings = 230,668,900,000 -  75,857,900,000 = 154,811,000,000

An expansionary fiscal policy will increase Aggregate Expenditure in the economy which in turn will increase National Income and Investment.


answered by: bluesky
Answer #2

By expenditure method of calculating National Income,

GDP \(=\) Consumption \((C)+\) Government Expenditure \((G)+\) Private Investment Expenditure(l) \(+\) Net Export (Export-Import)

GDP (in billion SAR) \(=222.526+140.199+1+0\)

\(517.536=222.526+140.199+1\)

\(\mathrm{I}=154.811\)

Sales \(\operatorname{Tax}=15 \%\) of \(222.526=33.3789\)

Tax on Investment \(=20 \%\) of \(154.811=30.9622\)

Total Tax \(=64.34110\) billion SAR

b) Investment has already been calculated. It is \(154.811\) billion SAR

c) Private Saving = Income - Taxes - Consumption

$$ \begin{array}{l} =517.536-64.34110-222.526 \\ =230.66890 \text { billion SAR } \end{array} $$

d) Public Saving = Taxes - Govt Expenditure

$$ =64.34110-140.199=-75.8579 \text { billion SAR } $$

e) National Saving = Private Saving \(+\) Public Saving

$$ =230.6689+(-75.8579)=154.8110 \text { billion SAR } $$

f) The government can decrease tax rates so that disposable income rises and hence, there can be a rise in savings of consumers inducing investment. It can also increase it's own expenditure to fuel aggregate demand by spending on capital investment.

2) The real money supply is measured in terms of Prices of commodities.

RealMoney \(=\frac{\text { NominalMoneySupply }}{\text { Price }}\)

So, if the price of a commodity rises, the purchasing power (with the nominal supply constant) falls. For example, if the Price is 10 and the nominal money supply is 60, we can purchase 6 units. But if the price rises to 12, then we can buy 5 units. Hence the value of money has fallen.

answered by: Dr.E